OpenStudy (maheshmeghwal9):

Write this in polar{trigonometrical form}: - \[-\sin 110^o+i \cos 110^o.\]

OpenStudy (anonymous):

evaluate is what you need to do

OpenStudy (maheshmeghwal9):

but i have done many questions but this is different

OpenStudy (anonymous):

actually, it is in polar form already

OpenStudy (anonymous):

weird that it says "write in trig from" when it is already in trig form

OpenStudy (maheshmeghwal9):

no; i tell u its answer but i don't know how can i gt that

OpenStudy (anonymous):

hold the phone. it says "write in trig form" and i am looking at trig functions right?

OpenStudy (maheshmeghwal9):

\[\cos (360^ok+200^o)+i \sin (360^ok+200^o).\]

OpenStudy (anonymous):

you could rewrite in standard from as \(a+bi\) or you could write as \(e^{i\theta}\) if you like

OpenStudy (anonymous):

ooooooooooooooooh i see !

OpenStudy (anonymous):

sorry my fault. they want you to write this as \(\cos(\theta)+i\sin(\theta)\)

OpenStudy (maheshmeghwal9):

np:) but now how to get it?

OpenStudy (maheshmeghwal9):

where the hell this 200 came from?

OpenStudy (anonymous):

ok lets see \(\cos(x)=\sin(90-x)\) so \(\cos(110)=\sin(90-110)=\sin(-20)\) no that doesn't do it, but it is a start

OpenStudy (maheshmeghwal9):

k

OpenStudy (anonymous):

that is the idea though, switch from cosine to sine and vice versa

OpenStudy (maheshmeghwal9):

ya u r right

OpenStudy (anonymous):

you got it from here?

OpenStudy (maheshmeghwal9):

no

OpenStudy (anonymous):

dumbass question if you don't mind me saying so

OpenStudy (maheshmeghwal9):

lol

OpenStudy (anonymous):

weird that i am getting a slightly different answer

OpenStudy (maheshmeghwal9):

Oh i gt it:) \[-\sin 110^o+i \cos 110^o.\]\[\implies - \cos (90-110)+i \sin (90-110)\]\[\implies - \cos (-20) + i \sin (-20)\]\[\implies \cos (180 - (-20))+ i \sin (180-(-20))= \cos (360^ok+200^o)+i \sin (360^ok+200^o)\]

OpenStudy (anonymous):

\(-\sin(110)=\sin(-110)\) because sine is odd. to switch to cosine we can use \(\cos(90-x)=\sin(x)\) so \(90+110=200\) ok got it

OpenStudy (anonymous):

oh and you got it too faster than i did

OpenStudy (anonymous):

still a dumbass question, has nothing at all to do with complex numbers, just changing cosines to sines etc

OpenStudy (maheshmeghwal9):

lol u helped me lot. thanx a lot:)